OpenStudy (anonymous):

In a plane there is a triangle B described by the properties: x<=2, y<=2, x+y>=2. Calculate the double integral. I can plot it but somehow I get a negative result.

OpenStudy (anonymous):

are we finding area or volume?

OpenStudy (anonymous):

volume

OpenStudy (anonymous):

then we need a function to integrate over given region

OpenStudy (anonymous):

oh, I'm sorry...I forgot half of the question

OpenStudy (anonymous):

the function that needs to be integrated is 2x-3y

OpenStudy (anonymous):

do you mean, what I calculated?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

\[x <= 2 \] \[y <= 2 \] \[x + y >= 2\] \[y >= -x + 2\]

OpenStudy (anonymous):

That's was wrong \[\int _0^2\int _{2-x}^22 x-3y dydx\]

OpenStudy (anonymous):

so the outer integral goes from 0 to 2 and the inner integral goes from 2-x to 2? because I got the same but still get a negative result

OpenStudy (anonymous):

did you get -8/3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then , you sir are correct

OpenStudy (anonymous):

but is this a realistic result...I mean after all it's a negative volume

OpenStudy (anonymous):

this is what we call as signed volume, yes that makes no sense. Remember in calc 2, we we integrate area under the x axis we get negative area; this is same thing

OpenStudy (anonymous):

ok, I think I understand it...plus I actually got my math right, so that's good....thanks a lot